# Interpretable Hybrid-Rule Temporal Point Processes

### v2

@@ Abstract
Temporal Point Processes (TPPs) are widely used for modeling event sequences in various medical domains, such as disease onset prediction, progression analysis, and clinical decision support. Although TPPs effectively capture temporal dynamics, their lack of interpretability remains a critical challenge. Recent advancements have introduced interpretable TPPs. However, these methods fail to incorporate numerical features, thereby limiting their ability to generate precise predictions. To address this issue, we propose Hybrid-Rule Temporal Point Processes (HRTPP), a novel framework that integrates temporal logic rules with numerical features, improving both interpretability and predictive accuracy in event modeling. HRTPP comprises three key components: basic intensity for intrinsic event likelihood, rule-based intensity for structured temporal dependencies, and numerical feature intensity for dynamic probability modulation. To effectively discover valid rules, we introduce a two-phase rule mining strategy with Bayesian optimization. To evaluate our method, we establish a multi-criteria assessment framework, incorporating rule validity, model fitting, and temporal predictive accuracy. Experimental results on real-world medical datasets demonstrate that HRTPP outperforms state-of-the-art interpretable TPPs in terms of predictive performance and clinical interpretability. In case studies, the rules extracted by HRTPP explain the disease progression, offering valuable contributions to medical diagnosis. The code is available at https://github.com/yy-c/HRTPP.
Index Terms:
Temporal Point Processes, Interpretability, Event Sequence Modeling, Logic Rules
I Introduction
@@
II Related Work
II-1 Temporal Point Process
@@
Temporal Point Process (TPP) models provide a probabilistic framework for capturing event sequences in continuous time. A key challenge in TPP research lies in designing intensity functions that are both expressive and interpretable. Efforts to improve interpretability in TPPs have led to the development of attention-based and rule-based approaches. Attention-based approaches such as SA-HP [18] and THP [12] incorporate self-attention mechanisms to capture long-range dependencies in event sequences. ITHP [19] further explains THP by algebraic operations. Although these methods improve flexibility, the trade-off between accuracy and interpretability remains an open problem. Rule-based TPP models focus on enhancing interpretability through structured rule sets. TLPP [20] first introduces logic rules to model event intensities. TELLER [14] introduces a principled approach to mine rules that explain temporal dependencies. Based on this, CLUSTER [4] and NS-TPP [21] design different deep neural network structures to accelerate rule mining. Clock Logic Neural Network (CLNN) [22] incorporates weighted clock logic formulas as rules to model event interactions. The recent trend in TPP research highlights the need to balance both explainable and effective aspects, integrating structured rule-based approaches with deep learning models for more explainable and effective event modeling.
II-2 Rule Mining
@@
Rule mining focuses on automatically discovering logical rules from event sequence data to explain event occurrence patterns. Traditional methods for rule discovery primarily identify frequent patterns, but struggle with capturing temporal dependencies. Apriori [23] efficiently extracts frequent event sets but fails to model event order. Sequential pattern mining techniques, such as CM-SPADE [24] and VGEN [25], attempt to incorporate temporal orders but face difficulties in handling fine-grained timestamps. To overcome these limitations, TELLER [14] formulates rule discovery as a maximum likelihood problem to generate high quality rules, but requires a lot of computation. CLUSTER [4] introduces an iterative framework based on the EM algorithm to improve rule mining efficiency. NS-TPP [21] represents predicates and logic rules as vector embeddings and introduces a neural-symbolic framework for rule mining. Clock Logic Neural Network (CLNN) [22] utilizes weighted clock logic formulas to introduce three novel rule forms and incorporate numerical values into the rules. HyperLogic [26] proposes the integration of if-then logic rules within neural network architectures.
III Background
@@
TPPs are widely used to interpret complex event sequences over time. For example, in ICU monitoring, physicians analyze the progression of symptoms and lab results to make decisions. Understanding the logic rules among medical events is crucial for accurate diagnosis and treatment. This section briefly reviews temporal point processes and temporal logic rules.
III-A Temporal Point Processes
@@
Temporal Point Process (TPP) is mathematically represented as a stochastic process $\{t_{i}\}_{i=1}^{N}$ where each $\{t_{i}\}$ denotes the timestamp of an event, and $N$ is the number of observed events within a given time window. The core component of TPPs is the conditional intensity function $\lambda(t)$, which defines the instantaneous rate of event occurrence at time $t$, given the history $\mathcal{H}_{t}=\{t_{1},t_{2},\dots,t_{i-1}\}$:
$\lambda(t|\mathcal{H}_{t})=\lim_{\Delta t\to 0}\frac{\mathbb{P}\big{(}N(t+\Delta t)-N(t)=1\mid\mathcal{H}_{t}\big{)}}{\Delta t}.$ (1)
@@
This intensity function governs the likelihood of an event happening at a particular time. The overall behavior of a TPP is captured by the joint distribution of event times. By modeling $\lambda(t|\mathcal{H}_{t})$, TPPs can describe a wide range of event patterns, including those influenced by past occurrences, external features, or domain-specific rules.
III-B Temporal Logic Rules
@@
Temporal logic rules consist of two parts: temporal predicates and relations. Define a set of temporal predicates as $\mathcal{X}=\{X_{1},X_{2},\dots,X_{n},Y\}$, where each predicates $X_{u}$ is an independent variable, $Y$ is a target variable. Define a set of temporal relations as $\mathcal{C}$, which define how the variables are related over time. Logic rules are in the form of ”if-then”. A generalized temporal logic rule $R$ can be expressed in the form:
@@
where $C_{i}\in\mathcal{C}$ represents a temporal relation, and $\psi_{1},\psi_{2}$ are expressions that can either be variables from $\mathcal{X}$ or recursively defined temporal rules, such as $X_{u}$ or $\psi_{u}$. The temporal relations between two temporal points can be divided into four types: and, before, equal, after [27]. Actually, ”$X_{1}$ before $X_{2}$” and ”$X_{2}$ after $X_{1}$” are equivalent, so the final set of temporal relations is $\mathcal{C}=\{and,before,equal\}$.
IV Methodology
This section introduces the framework of Hybrid-Rule Temporal Point Processes (HRTPP), shown in Fig. 1. Given an event sequence $\mathcal{S}=\{(t_{i},k_{i},v_{i})\}_{i=1}^{N}$ consisting of $N$ events, each event type $k_{i}\in\{1,2,\dots,K\}$ is associated with an independent variable $X_{k}$ and a numerical feature $v_{i}$, providing contextual information. For instance, in ICU monitoring, an event $k_{i}$ ”elevated heart rate” at $t_{i}=2$ hours may have $v_{i}=120$ bpm. The conditional intensity function $\lambda(t|\mathcal{H}_{t})$ is extended to incorporate temporal logic rules alongside numerical attributes, formulated as follows:
@@ IV Methodology
IV-A Rule-Based Feature Construction
@@
IV-A1 Definitions of Temporal Rules
@@
Given two predicates $X_{u}$ and $X_{v}$ with timestamps $t_{u}$ and $t_{v}$, along with a target event $Y$ occurring at $t_{y}>\max(t_{u},t_{v})$, the three types of temporal rules are mathematically defined as follows. Before, where $X_{u}$ occurs before $X_{v}$ within a small time tolerance $\delta$, formulated as $\mathcal{C}_{before}(X_{u},X_{v})=\mathbb{I}(t_{u}-t_{v}<-\delta)$. Equal, where $X_{u}$ and $X_{v}$ occur approximately at the same time, formulated as $\mathcal{C}_{equal}(X_{u},X_{v})=\mathbb{I}(|t_{u}-t_{v}|\leq\delta)$. And, which simply states that both events occur without temporal constraints, formulated as $\mathcal{C}_{and}(X_{u},X_{v})=1$. Here, $\mathbb{I}(\cdot)$ represents the indicator function. These temporal rules provide a structured mechanism for capturing event dependencies.
IV-A2 Rule Encoder
To effectively incorporate temporal rules into the TPP framework, we encode hybrid-rule information into a continuous vector space. Based on the definitions of temporal rules, we construct hybrid-rule event sequences. For instance, consider the temporal rule ”$X_{u}$ before $X_{v}$to $Y$”. The corresponding hybrid-rule event sequence records the timestamps when the condition $X_{u}$ before $X_{v}$ is satisfied. Specifically, this sequence tracks each moment where the constraint $\mathcal{C}_{before}(X_{u},X_{v})$ transitions from 0 to 1. The encoding function for a rule $R_{j}$ is defined as:
@@ IV-A2 Rule Encoder
IV-A3 Numerical Feature Encoder
Numerical features provides quantitative insight into the condition of a patient. Numerical feature encoder captures the dynamic influence of numerical attributes over time. Given an event sequence $\mathcal{S}=\{(t_{i},k_{i},v_{i})\}_{i=1}^{N}$, the sequence corresponding to predicate $X_{k}$ is expressed as $\{(t_{i},k_{i},v_{i})\}_{k_{i}=k}$. However, not all predicates directly influence the target event. To address this, we introduce a masking mechanism to filter out irrelevant predicates. The mask is defined in accordance with the given rule set, where only predicates appearing in rules associated with the target event remain active. Formally, the mask set is denoted as =$\mathcal{M}=\{m_{k}\}_{k=1}^{K}$, and the valid predicate set $\mathcal{X}_{V}\subset\mathcal{X}$ is defined as the collection of all variables appearing in the rules. The mask function is defined as $m_{k}=1$ if $X_{k}\in\mathcal{X}_{V}$, and $m_{k}=0$ otherwise. Using this masking mechanism and the sequence $\{(t_{i},k_{i},v_{i})\}_{k_{i}=k}$, we define the numerical feature encoding function for predicate $X_{k}$ as:
@@ IV-A3 Numerical Feature Encoder
IV-B Hybrid-Rule Temporal Point Processes
@@ IV-B1 Basic Intensity
IV-B2 Rule-Based Intensity
@@
IV-B3 Numerical Feature Intensity
@@
where $\Theta=\{\boldsymbol{\alpha},\boldsymbol{\beta},\gamma,\lambda_{0}\}$ is learnable parameters set, $\boldsymbol{\alpha}=\{\alpha_{j}\}_{j=1}^{|\mathcal{R}|}$ is the set of weights associated with temporal rules, $\boldsymbol{\beta}=\{\beta_{i}\}_{i=1}^{K}$ is the set of weights associated with numerical features. This overall intensity function serves as the foundation for modeling event occurrences in HRTPP framework, enabling robust modeling in event prediction and rule mining.
IV-C Training and Prediction
@@
IV-C1 Loss Function
[…]
@@ IV-C2 Event Prediction
[…]
@@
IV-D Rule Mining and Optimization
@@
IV-D1 Rule Candidate Generation
Temporal rules are derived from historical event sequences to construct a structured search space. Without constraints, the combination of predicates and relations can be excessively large. To address this, rule length is restricted, and predicate filtering is applied. Firstly, rule length significantly impacts interpretability. Overly complex rules hinder comprehension, whereas single-predicate rules lack temporal expressiveness. Based on medical guidelines, a maximum of two or three predicates is imposed. Secondly, predicate filtering is conducted via pre-training using a rule-free TPP model as a baseline. Each predicate is evaluated individually, and those reducing model loss are retained. The resulting filtered predicates are used to generate candidate rules, ensuring that each rule contains at least one valid predicate. This process refines the search space, facilitating efficient rule selection.
IV-D2 Rule Optimization via Bayesian Optimization
@@
where $\log\mathcal{L}(\mathcal{R};\Theta)$ denotes the model likelihood given the rule set $\mathcal{R}$. To maintain efficiency and interpretability, the rule set size $|\mathcal{R}|$ is constrained.
@@
V Experiments
V-A Experimental Setup
Datasets. We use four real-world disease-specific datasets: AKI (Acute Kidney Injury), Stroke, Sepsis, and CAD (Coronary Artery Disease). These medical datasets are extracted from MIMIC-IV [28], a large-scale medical database that contains electronic health records. Table I presents the statistical information of these datasets. As shown in the table, these four datasets increase in size from top to bottom. In the AKI dataset, disease progression to Phase III serves as the target event. The Stroke dataset focuses on patients transitioning from moderate to severe conditions. In the Sepsis dataset, the occurrence of low urine output is tracked as a target of organ failure. The CAD dataset defines patient mortality as the target event.
Baselines. We choose three state-of-the-art TPP models that are capable of rule mining.
TEmporal Logic rule LearnER (TELLER) [14]: This model can discover temporal logic rules by rule-extension-first-search or rule-addition-first-search.
Clock Logic Neural Network (CLNN) [22]: This model introduces a time-aware framework to provide weighted clock logic formulas as rules.
CLUSTER [4]: This model automatically extracts ”if-then” logic rules to interpret observational events.
Evaluation Metrics. To assess the performance of the models, we conducted comprehensive experiments. Each dataset is split into an 80% training set and a 20% test set. The evaluation metrics used to measure predictive performance include Negative Log-Likelihood (NLL), Root Mean Squared Error (RMSE) and Rule Accuracy (Acc). Rule accuracy means the proportion of correct rules.
TABLE I: Datasets statistics.
Dataset Sequences Types Events Seq. Length Avg. Length

### v3

@@ Abstract
Temporal Point Processes (TPPs) are widely used for modeling event sequences in various medical domains, such as disease onset prediction, progression analysis, and clinical decision support. Although TPPs effectively capture temporal dynamics, their lack of interpretability remains a critical challenge. Recent advancements have introduced interpretable TPPs. However, these methods fail to incorporate numerical features, thereby limiting their ability to generate precise predictions. To address this issue, we propose Hybrid-Rule Temporal Point Processes (HRTPP), a novel framework that integrates temporal logic rules with numerical features, improving both interpretability and predictive accuracy in event modeling. HRTPP comprises three key components: basic intensity for intrinsic event likelihood, rule-based intensity for structured temporal dependencies, and numerical feature intensity for dynamic probability modulation. To effectively discover valid rules, we introduce a two-phase rule mining strategy with Bayesian optimization. To evaluate our method, we establish a multi-criteria assessment framework, incorporating rule validity, model fitting, and temporal predictive accuracy. Experimental results on real-world medical datasets demonstrate that HRTPP outperforms state-of-the-art interpretable TPPs in terms of predictive performance and clinical interpretability. In case studies, the rules extracted by HRTPP explain the disease progression, offering valuable contributions to medical diagnosis. The supplementary material is available at https://github.com/yy-c/HRTPP.
* These authors contributed equally to this work.
1 Introduction
@@
2 Related Work
2.0.1 Temporal Point Processes.
@@ 2 Related Work
Temporal Point Process (TPP) models provide a probabilistic framework for capturing event sequences in continuous time. A key challenge in TPP research lies in designing intensity functions that are both expressive and interpretable. Efforts to improve interpretability in TPPs have led to the development of attention-based and rule-based approaches. Attention-based approaches such as SA-HP [28] and THP [29] incorporate self-attention mechanisms to capture long-range dependencies in event sequences. ITHP [19] further explains THP by algebraic operations. Although these methods improve flexibility, the trade-off between accuracy and interpretability remains an open problem. Rule-based TPP models focus on enhancing interpretability through structured rule sets. TLPP [16] first introduces logic rules to model event intensities and defines temporal rules as a combination of temporal predicates and relations. TELLER [15] introduces a principled approach to mine rules that explain temporal dependencies. Based on this, CLUSTER [14] and NS-TPP [27] design different deep neural network structures to accelerate rule mining. Clock Logic Neural Network (CLNN) [24] incorporates weighted clock logic formulas as rules to model event interactions. The recent trend in TPP research highlights the need to balance both explainable and effective aspects, integrating structured rule-based approaches with deep learning models for more explainable and effective event modeling.
2.0.2 Rule Mining.
@@
Rule mining focuses on automatically discovering logical rules from event sequence data to explain event occurrence patterns. Traditional methods for rule discovery primarily identify frequent patterns, but struggle with capturing temporal dependencies. Apriori [1] efficiently extracts frequent event sets but fails to model event order. Sequential pattern mining techniques, such as CM-SPADE [7] and VGEN [8], attempt to incorporate temporal orders but face difficulties in handling fine-grained timestamps. To overcome these limitations, TELLER [15] formulates rule discovery as a maximum likelihood problem to generate high quality rules, but requires a lot of computation. CLUSTER [14] introduces an iterative framework based on the Expectation-Maximization algorithm to improve rule mining efficiency. NS-TPP [27] represents predicates and logic rules as vector embeddings and introduces a neural-symbolic framework for rule mining. Clock Logic Neural Network (CLNN) [24] utilizes weighted clock logic formulas to introduce three novel rule forms and incorporate numerical values into the rules. HyperLogic [26] proposes the integration of if-then logic rules within neural network architectures.
3 Background
@@
TPPs are widely used to interpret complex event sequences over time. For example, in intensive care unit (ICU) monitoring, physicians analyze the progression of symptoms and lab results to make decisions. Understanding the logic rules among medical events is crucial for accurate diagnosis and treatment. This section briefly reviews temporal point processes and temporal logic rules.
3.1 Temporal Point Processes
@@
Temporal Point Process (TPP) is mathematically defined as a stochastic process $\{t_{i}\}_{i=1}^{N}$ where each $t_{i}$ denotes the timestamp of an event, and $N$ is the number of observed events. The core component of TPPs is the conditional intensity function $\lambda(t)$, which defines the instantaneous rate of event occurrence at time $t$, conditioned on the history $\mathcal{H}_{t}=\{t_{1},t_{2},\dots,t_{i}\}$ up to but not including $t$:
$\lambda(t|\mathcal{H}_{t})=\lim_{\Delta t\to 0}\frac{\mathbb{P}\left(N(t+\Delta t)-N(t)=1\mid\mathcal{H}_{t}\right)}{\Delta t}.$ (1)
@@
Here, $\Delta t$ is a small time interval, and $N(t)$ denotes the counting process up to $t$. This intensity function governs the likelihood of an event happening at a particular time. The overall dynamics are characterized by the joint distribution. By modeling $\lambda(t|\mathcal{H}_{t})$, TPPs can describe a wide range of event patterns, including those influenced by past occurrences, external features, or domain-specific rules.
3.2 Temporal Logic Rules
@@
Temporal logic rules consist of two parts: temporal predicates and relations. Define a set of predicates as $\mathcal{X}=\{X_{1},X_{2},\dots,X_{n},Y\}$, where each predicates $X_{u}$ is an independent variable, $Y$ is a target variable. Define a set of temporal relations as $\mathcal{C}$, which define how the variables are related over time. Rules are in the form of "if-then". A generalized temporal logic rule $R$ can be expressed as:
@@
where $C_{i}\in\mathcal{C}$ represents a temporal relation, and $\psi_{1},\psi_{2}$ are expressions that can either be variables from $\mathcal{X}$ or recursively defined temporal rules. For example, $\psi$ can represent the variable "high heart rate" or the rule "high heart rate before hypertension". The temporal relations between two temporal points can be divided into four types: and, before, equal, after [2]. Actually, $X_{1}$ before $X_{2}$" and $X_{2}$ after $X_{1}$" are equivalent, so the final set of temporal relations is $\mathcal{C}=\{and,before,equal\}$.
4 Methodology
This section introduces the framework of Hybrid-Rule Temporal Point Processes (HRTPP), shown in Fig. 1. Given an event sequence $\mathcal{S}=\{(t_{i},k_{i},v_{i})\}_{i=1}^{N}$ consisting of $N$ events, each event type $k_{i}\in\{1,2,\dots,K\}$ is associated with an independent variable $X_{k}$ and a numerical feature $v_{i}$, providing contextual information. For instance, in ICU monitoring, an event $k_{i}$ "elevated heart rate" at $t_{i}=2$ hours may have $v_{i}=120$ bpm. The conditional intensity function $\lambda(t|\mathcal{H}_{t})$ is extended to incorporate temporal logic rules alongside numerical attributes, formulated as follows:
@@ 4 Methodology
4.1 Rule-Based Feature Construction
@@
4.1.1 Definitions of Temporal Rules.
@@
Given two predicates $X_{u}$ and $X_{v}$ with timestamps $t_{u}$ and $t_{v}$, along with a target event $Y$ occurring at $t_{y}>\max(t_{u},t_{v})$ and the indicator function $\mathbb{I}(\cdot)$, the three types of temporal rules are mathematically defined as follows. Before, where $X_{u}$ occurs before $X_{v}$ within a small time tolerance $\delta$, formulated as $\mathcal{C}_{before}(X_{u},X_{v})=\mathbb{I}(t_{u}-t_{v}<-\delta)$. Equal, where $X_{u}$ and $X_{v}$ occur approximately at the same time, formulated as $\mathcal{C}_{equal}(X_{u},X_{v})=\mathbb{I}(|t_{u}-t_{v}|\leq\delta)$. And, which simply states that both events occur without temporal constraints, formulated as $\mathcal{C}_{and}(X_{u},X_{v})=1$. These temporal rules provide a structured mechanism for capturing event dependencies.
4.1.2 Rule Encoder.
To effectively incorporate temporal rules into the TPP framework, we encode hybrid-rule information into a continuous vector space. Based on the definitions of temporal rules, we construct hybrid-rule event sequences. For instance, consider the temporal rule $X_{u}$ before $X_{v}$ $\to$ YY". The corresponding hybrid-rule event sequence records the timestamps when the condition $X_{u}$ before $X_{v}$ is satisfied. Specifically, this sequence tracks each moment where the constraint $\mathcal{C}_{before}(X_{u},X_{v})$ transitions from 0 to 1. The encoding function for a rule $R_{j}$ is defined as:
@@
4.1.3 Numerical Feature Encoder.
Numerical features provide quantitative insight into the condition of a patient. The numerical feature encoder captures the dynamic influence of numerical attributes over time. Given an event sequence $\mathcal{S}=\{(t_{i},k_{i},v_{i})\}_{i=1}^{N}$, the sequence corresponding to the predicate $X_{k}$ is expressed as $\{(t_{i},k_{i},v_{i})\}_{k_{i}=k}$. Generally, not all predicates directly influence the target event. Therefore, we introduce a masking mechanism to filter out irrelevant predicates. The mask is defined according to the given rule set, where only the predicates appearing in rules associated with the target event remain active, ensuring that the retained variables are indeed the relevant ones. Formally, the mask set is denoted as $\mathcal{M}=\{m_{k}\}_{k=1}^{K}$, and the valid predicate set $\mathcal{X}_{V}\subset\mathcal{X}$ denotes the collection of all variables appearing in the rules. The mask function is defined as $m_{k}=1$ if $X_{k}\in\mathcal{X}_{V}$, and $m_{k}=0$ otherwise. Using this masking mechanism and the sequence $\{(t_{i},k_{i},v_{i})\}_{k_{i}=k}$, we define the numerical feature encoding function for predicate $X_{k}$ as:
@@
4.2 Hybrid-Rule Temporal Point Processes
@@
4.2.2 Rule-Based Intensity.
@@
4.2.3 Numerical Feature Intensity.
@@
where $\boldsymbol{\alpha}=\{\alpha_{j}\}_{j=1}^{|\mathcal{R}|}$ is the set of weights associated with temporal rules, $\boldsymbol{\beta}=\{\beta_{i}\}_{i=1}^{K}$ is the set of weights associated with numerical features. The learnable parameters set is denoted as $\Theta=\{\boldsymbol{\alpha},\boldsymbol{\beta},\gamma,\lambda_{0}\}$. This overall intensity function serves as the foundation for modeling event occurrences in HRTPP framework, enabling robust modeling in event prediction and rule mining.
4.3 Training and Prediction
@@
4.3.1 Loss Function.
The HRTPP model is trained by maximizing the likelihood of observed event sequences within the temporal point process framework. Given an event sequence $\mathcal{S}={(t_{i},k_{i},v_{i})}_{i=1}^{N}$ and a rule set $\mathcal{R}$, the likelihood function is formulated as $\mathcal{L}(\Theta)=\prod_{i=1}^{N}\lambda(t_{i}|\mathcal{H}_{t_{i}})\exp\left(-\int_{0}^{t_{N}}\lambda(t|\mathcal{H}_{t})dt\right)$, where $\Theta$ is the set of learnable model parameters. To facilitate optimization, we minimize the negative log-likelihood (NLL) as the loss function:
$\mathcal{L}_{\text{NLL}}(\Theta)=-\sum_{i=1}^{N}\log\lambda(t_{i}|\mathcal{H}_{t_{i}})+\int_{0}^{t_{N}}\lambda(t|\mathcal{H}_{t})dt.$ (9)
@@
$p(\hat{t}|\mathcal{H}_{t})=\lambda(\hat{t}|\mathcal{H}_{t})\exp\left(-\int_{t}^{\hat{t}}\lambda(s|\mathcal{H}_{s})ds\right).$ (10)
@@
4.4 Rule Mining and Optimization
@@
4.4.1 Rule Candidate Generation.
Temporal rules are derived from historical event sequences to construct a structured search space. Without constraints, the combination of predicates and relations can be excessively large. For instance, any rule containing $m$ predicates interconnected by (m1)(m-1) relations yields a combinatorial space of $K^{m}C^{m-1}$ possible configurations, with $K$ and $C$ being the total number of the predicates and relations respectively. To address this, rule length is restricted, and predicate filtering is applied. Firstly, rule length significantly impacts interpretability. Overly complex rules hinder comprehension, whereas single-predicate rules lack temporal expressiveness. Based on medical guidelines, a maximum of two or three predicates is imposed. Secondly, predicate filtering is conducted via pre-training using a rule-free TPP model as a baseline. Each predicate is evaluated individually, and those reducing model loss are retained. The resulting filtered predicates are used to generate candidate rules, ensuring that each rule contains at least one valid predicate. This process refines the search space, facilitating efficient rule selection.
4.4.2 Rule Optimization via Bayesian Optimization.
@@
where $\log\mathcal{L}(\mathcal{R};\Theta)$ denotes the model likelihood given the rule set $\mathcal{R}$. To maintain efficiency and interpretability, the rule set size $|\mathcal{R}|$ is is constrained to a fixed constant.
@@
5 Experiments
5.1 Experimental Setup
Datasets. We use four real-world disease-specific datasets: AKI (Acute Kidney Injury), Stroke, Sepsis, and CAD (Coronary Artery Disease). These medical datasets are extracted from MIMIC-IV [12], a large-scale medical database that contains electronic health records. Table 1 presents the statistical information of these datasets. The number sign # denotes a summation operation. As shown in the table, these four datasets increase in size from top to bottom. In the AKI dataset, disease progression to Phase III serves as the target event. The Stroke dataset focuses on patients transitioning from moderate to severe conditions. In the Sepsis dataset, the occurrence of low urine output is tracked as a target of organ failure. In the CAD dataset, patient mortality is defined as the target event.
Baselines. We adopted three state-of-the-art models with explicit rule-mining capabilities: TEmporal Logic rule LearnER (TELLER) [15], Clock Logic Neural Network (CLNN) [24], and CLUSTER [14].
Evaluation Metrics. To assess the performance of the models, we conducted comprehensive experiments. Each dataset is randomly split into an 80% training set and a 20% test set. The evaluation metrics used to measure predictive performance include Negative Log-Likelihood (NLL), Root Mean Squared Error (RMSE) and Rule Accuracy (Acc). Rule accuracy means the proportion of correct rules.
Table 1: Datasets statistics.
Dataset # Sequences # Types # Events Seq. Length Avg. Length